What is -2/3,0.6,3/4,-7/4 and -0.3 ordered from least to greatest?
Answer by stanbon(75887)

What is
-2/3 = -0.667...
0.6 = 0.600..
3/4 = 0.750
-7/4 = -1.75
-0.3 = -0.300
ordered from least to greatest?
-1.75 ; -0.667 ; -0.300 ; 0.600 ; 0.750
====================
Cheers,
Stan H.
====================